OpenStudy (anonymous):

What is the solution to 4x -14 + 6 = 12?

OpenStudy (mathstudent55):

First add the numbers on the left side together. Can you do that?

OpenStudy (anonymous):

They equal -4?

OpenStudy (mathstudent55):

By numbers I mean the plain numbers that have no variable, -14 and 6. You can't add numbers with variables to numbers without variable.

OpenStudy (mathstudent55):

The 4x remains as 4x. The -14 and 6 are added together.

OpenStudy (anonymous):

-8?

OpenStudy (mathstudent55):

Good, so now you have: 4x - 8 = 12, ok?

OpenStudy (anonymous):

I'm with you so far.

OpenStudy (anonymous):

4x - 14 + 6 =12 4x - 8 - 12 = 0 4x - 20 = 0 4x = 20 x = 5

OpenStudy (mathstudent55):

The solution to your equation is x = some number. We want x by itself, but there are two numbers still on the left side, 4 multiplying x and -8 being subtracted. We need to get rid of both numbers from the left side. We have to do it in a way that math allows us to do it in.

OpenStudy (mathstudent55):

We are allowed to add the same number to both sides of an equation. Since there is an 8 being subtracted on the left side, if we add 8 to the left side, the -8 will be gone. But whatever we add to the left side, we must also do to the right side. So the next step is to add 8 to both sides.

OpenStudy (mathstudent55):

4x - 8 = 12 +8 +8 -----------------Add 4x = 20

OpenStudy (anonymous):

20 divided by 4 = 5. hence, x = 5

OpenStudy (anonymous):

I get it now, thank you.

OpenStudy (mathstudent55):

Now we got rid of the -8. We only have one more step to do. The x is being multiplied by 4. We need to get rid of it from the left side. We can divide both sides of an equation by the same number. Since x is being multiplied by 4, we need to divide the left side by 4, since 4/4 = 1. We must divide the right side by 4 also. 4x/4 = 20/4 x = 5

OpenStudy (mathstudent55):

The final answer is x = 5